One year ago Sam purchased bonds for $100,000. He just sold them for $120,000. During the year the price level rose by 5%. If the tax rate on capital gains is 20%, how much did Sam gain in real terms?
Answer #1

Capital Gain=Sale price-Purchase price=120000-100000=$20000

Capital gain tax=20000*20%=$4000

After tax sale value=120000-4000=$116,000

Value of $116,000 in real terms=116000/(1+5%)=$110476.19

Real gain in absolute terms=110476.19-100000=$10476.19

Real Gain=(110476.19-100000)/100000=10.48%